# Mambular: A Sequential Model for Tabular Deep Learning

### v1

@@ Abstract
The analysis of tabular data has traditionally been dominated by gradient-boosted decision trees (GBDTs), known for their proficiency with mixed categorical and numerical features. However, recent deep learning innovations are challenging this dominance. We introduce Mambular, an adaptation of the Mamba architecture optimized for tabular data. We extensively benchmark Mambular against state-of-the-art models, including neural networks and tree-based methods, and demonstrate its competitive performance across diverse datasets. Additionally, we explore various adaptations of Mambular to understand its effectiveness for tabular data. We investigate different pooling strategies, feature interaction mechanisms, and bi-directional processing. Our analysis shows that interpreting features as a sequence and passing them through Mamba layers results in surprisingly performant models. The results highlight Mambular’s potential as a versatile and powerful architecture for tabular data analysis, expanding the scope of deep learning applications in this domain. The source code is available at https://github.com/basf/mamba-tabular.
@@ 1 Introduction
Gradient-boosted decision trees (GBDTs) have historically dominated the field of tabular data analysis, with widely used variants including XGBoost, LightGBM, and CatBoost. These models excel in managing the mix of categorical and numerical features typical of tabular data (Grinsztajn et al.,, 2022). Historically, deep learning models struggled with tabular data due to its inherent complexity and diversity, often failing to outperform GBDTs due to challenges in handling missing values, diverse feature types, and the necessity for extensive preprocessing (Borisov et al.,, 2022). However, recent advancements in deep learning are gradually challenging this paradigm by introducing innovative architectures that leverage advanced mechanisms to capture complex feature dependencies, promising significant improvements (Popov et al.,, 2019; Hollmann et al.,, 2022; Gorishniy et al.,, 2021).
@@
One of the most effective advancements in tabular deep learning is the application of attention mechanisms in models like TabTransformer (Huang et al.,, 2020), FT-Transformer (Gorishniy et al.,, 2021) and many more (Wang and Sun,, 2022; Thielmann et al., 2024b, ; Arik and Pfister,, 2021). These models leverage the attention mechanism to capture dependencies between features, offering a significant improvement over traditional approaches. FT-Transformers, in particular, have demonstrated robust performance across various tabular datasets, often surpassing the accuracy of GBDTs (McElfresh et al.,, 2024).
Additionally, more traditional models like Multi-Layer Perceptrons (MLPs) and ResNets have demonstrated improvements when well-designed and when the data undergoes thorough preprocessing (Gorishniy et al.,, 2021, 2022). These models have benefited especially from innovations in advanced preprocessing methods that make them more competitive.
More recently, the Mamba architecture (Gu and Dao,, 2023) has shown promising results in textual problems. Tasks previously dominated by Transformer architectures, such as DNA modeling and large language models (LLMs), have seen improvements with the application of Mamba models (Gu and Dao,, 2023; Schiff et al.,, 2024; Zhao et al.,, 2024).
@@
Several adaptations have demonstrated its versatility, such as Vision Mamba for image classification (Xu et al.,, 2024), video analysis (Yang et al.,, 2024; Yue and Li,, 2024) and point cloud analysis (Zhang et al.,, 2024; Liu et al.,, 2024). Additionally, the architecture has been adapted for time series problems, with notable successes reported by Patro and Agneeswaran, (2024), Wang et al., (2024) and Ahamed and Cheng, 2024b . Mamba has also been integrated into graph learning (Behrouz and Hashemi,, 2024) and imitation learning (Correia and Alexandre,, 2024). Further advancements have improved the language model, by e.g. incorporating attention (Lieber et al.,, 2024), Mixture of Experts (Pióro et al.,, 2024) or bi-directional sequence processing (Liang et al.,, 2024).
@@
These advancements underscore Mamba’s broad applicability, making it a powerful and flexible architecture for diverse tasks and data types. Similarly to the transformer architecture, the question arises whether the Mamba architecture can also be leveraged for tabular problems. While Ahamed and Cheng, 2024a has already demonstrated that the Mamba architecture can be used for tabular data, a deeper analysis regarding the model architecture as well as insights into optimizations for tabular datasets is needed.
@@
    We present Mambular, a tabular adaptation of Mamba, and showcase the applicability of sequential models to tabular problems.
@@
    Mambular is extensively benchmarked against several other competitive neural as well as tree-based methods, showcasing that a default Mambular model performs on par with or better than tree-based models across a wide range of datasets.
@@
    We analyze the effects of bi-directional processing as well as feature interaction layers on Mambular’s performance and compare classical pooling methods.
@@
    Lastly, we conduct a thorough analysis of the sequential nature of Mambular, exploring the effects of feature orderings in a sequential tabular model.
@@ 2 Methodology
For a tabular problem, let $\mathcal{D}=\{(\bm{x}^{(i)},y^{(i)})\}_{i=1}^{n}$ be the training dataset of size $n$ and let $y$ denote the target variable that can be arbitrarily distributed. Each input $\bm{x}=(x_{1},x_{2},\dots,x_{J})$ contains $J$ features (variables). Categorical and numerical features are distinguished as $\bm{x}\equiv(\bm{x}_{cat},\bm{x}_{num})$, with the complete feature vector denoted as $\bm{x}$. Further, let $x_{j(cat)}^{(i)}$ denote the $j$-th categorical feature of the $i$-th observation, and hence $x_{j(num)}^{(i)}$ denote the $j$-th numerical feature of the $i$-th observation.
The core of the Mambular architecture is a combination of the FT-Transformer (Gorishniy et al.,, 2021) and Mamba (Gu and Dao,, 2023). Following classical tabular transformer architectures, the categorical features are first encoded and embedded. In contrast to classical language models, each categorical feature has its own, distinct vocabulary to avoid problems with binary or integer encoded variables. Including $<$unk$>$ tokens additionally allows to easily deal with unknown or missing categorical values during training or inference.
Numerical features are mapped to the embedding space via a simple linear layer. However, since a single linear layer does not add information beyond a linear transformation, Periodic Linear Encodings, as introduced by Gorishniy et al., (2022), are used for all numerical features. Thus, each numerical feature is encoded before being passed through the linear layer for rescaling. Simple decision trees are used for detecting the bin boundaries, and depending on the task, either classification or regression is employed for the target-dependent encoding function $h_{j}(\bm{x}_{j(num)},y)$. Let $b_{t}$ denote the decision boundaries from the decision trees, and the encoding function is given in Eq. 1.
@@
[…]
Subsequently, the embeddings are passed jointly through a stack of Mamba layers. These include one-dimensional convolutional layers as well as a state-space (SSM) model (Gu et al.,, 2021; Hamilton,, 1994) . The feature matrix before being passed through the SSM model has a shape of (Batch Size) $\times$ J $\times$ (Embedding Dimension), later referenced as $N\times J\times d$. Importantly, the sequence length in a tabular context refers to the number of variables, and hence the second dimension, $J$, corresponds to the number of features rather than to the length of, e.g., a document. Given matrices:
$\mathbf{A}\in\mathbb{R}^{1\times 1\times d\times\delta},\quad\mathbf{B}\in\mathbb{R}^{N\times J\times 1\times\delta},\quad\Delta\in\mathbb{R}^{N\times J\times d\times 1},\bar{\mathbf{z}}\in\mathbb{R}^{N\times J\times d\times 1},$
@@
$\mathbf{h}_{j}=\exp\left(\Delta\odot_{3}\mathbf{A}\right)_{:,j,:,:}\odot_{1,2,3}\mathbf{h}_{j-1}+\left(\left(\Delta\odot_{1,2}\mathbf{B}\right)\odot_{1,2,3}\bar{\mathbf{z}}\right)_{:,j,:,:}.$ (2)
@@
In contrast to FT-Transformer (Gorishniy et al.,, 2021) and TabTransformer (Huang et al.,, 2020) Mambular truly iterates through all variables as if they are a sequence. Hence, feature interactions are detected sequentially; and the effect of feature position in a sequence is analyzed with respect to performance in section 4. Furthermore, it should be noted that in contrast to TabPFN (Hollmann et al.,, 2022), Mambular does not transpose dimensions and iterates over observations. Hence, training on large datasets is possible and it can scale well to any training data size, just as Mamba (Gu and Dao,, 2023) does.
@@
Before it is passed to the final task specific model head, several pooling techniques along the sequence axis are possible and tested, e.g. sum pooling, average pooling or simple stacking.
We conduct several experiments with [cls] tokens, and analyze the model’s performance when only passing the [cls] token to the task-specific head. Given the recurrent update of the hidden states when iterating over the variables, variable positioning was also analyzed, since the order of the variables could affect this sequential setup. The model is trained end-to-end by minimizing the task-specific loss, e.g., mean squared error for regression or categorical cross entropy for classification tasks. An overview of a forward pass in the model is given in Figure 2.
@@ 3 Experiments
Mambular is benchmarked across a variety of datasets against the currently top-performing models (McElfresh et al.,, 2024). FT-Transformer (Gorishniy et al.,, 2021), TabTransformer (Huang et al.,, 2020), the consistently well-performing XGBoost (Grinsztajn et al.,, 2022; McElfresh et al.,, 2024), a baseline Multi-Layer Perceptron, and a ResNet are the models which contributed in this benchmark. Note, that TabPFN (Hollmann et al.,, 2022) is not included because it is not suited for larger datasets.
@@
We perform 5-fold cross-validation for all datasets and report average results as well as standard deviations. For all neural models, PLE encodings (Eq. 1) with a maximum number of bins equal to the model dimension are used (128 for most models, including MLP and ResNet). All categorical features are integer-encoded. For the regression tasks, the targets are normalized. Mean Squared Error (MSE) and Area Under the Curve (AUC) statistics are reported for regression and classification tasks, respectively. TabTransformer, FT-Transformer, and Mambular use the same architectures for embeddings as well as for the task-specific head, which consists of a single output layer without activation or dropout. For FT-Transformer, the [cls] token embedding for final prediction is used since it has been shown to improve performance (Thielmann et al., 2024b, ; Gorishniy et al.,, 2021).
@@
Several shared parameters are used for all neural models: a starting learning rate of 1e-04, weight decay of 1e-06, an early stopping patience of 15 epochs with respect to the validation loss, a maximum of 200 epochs for training, and a small learning rate decay with a factor of 0.1 with a patience of 10 epochs with respect to the validation loss. Additionally, a universal batch size of 128 is used, and the best model with respect to the validation loss is returned for testing. For TabTransformer, FT-Transformer, and Mambular, the same embedding functions are used. For the benchmarks, a vanilla Mambular architecture is used, utilizing average pooling, no feature interaction layer, and no bi-directional processing. Additionally, the columns/sequence are always sorted with numerical features first, followed by categorical features. Within these two groups, the features are sorted as they were originally provided in the dataset from the UCI Machine Learning Repository. Details on the used datasets and preprocessing can be found in Appendix A. Details on the model architectures and hyperparameters can be found in Appendix C.
MambaTab
In addition to these popular tabular models, we tested out the architecture proposed by Ahamed and Cheng, 2024a . MambaTab is the first architecture to leverage Mamba blocks for tabular problems. However, the authors propose using a combined linear layer to project all inputs into a single feature representation, transforming the features into a pseudo-sequence of fixed length 1. This approach simplifies the recursive update from Eq. 2 into a matrix multiplication and makes the model resemble a ResNet due to the residual connections in the final processing. Utilizing a sequential model with a sequence length of 1 does not fully exploit the strengths of sequential processing, as it reduces the model’s capacity to capture dependencies across multiple features.
We tested the architecture proposed by Ahamed and Cheng, 2024a and could achieve similar results for shared datasets, but overall found MambaTab to perform similar to a ResNet, aligning with expectations (see Table 2 and 23). Additionally, we experimented with transposing the axes to create an input matrix of shape (1) $\times$ (Batch Size) $\times$ (Embedding Dimension), as outlined in their implementation. While this approach draws on ideas from TabPFN (Hollmann et al.,, 2022), it did not lead to performance improvements in our experiments. When using PLE encodings and increasing the number of layers and dimensions compared to the default implementation from Ahamed and Cheng, 2024a we are able to increase performance. See Appendix C for a further discussion on MambaTab.
Comparison to XGBoost
Comparing mambular to XGBoost, we find that with default hyperparameter settings, Mambular performs as good, if not minimally better than XGBoost. On 12 datasets it outperforms XGBoost significantly for 4 datasets at the 10% significance level, while XGboost outperforms Mambular for 2 datasets at the 10 % significance level. The $p$-values from simple t-tests over the folds are reported for each dataset and the testing is oriented on Gorishniy et al., (2021). When adjusting for multiple testing via Benjamini-Hochberg (Ferreira and Zwinderman,, 2006; Benjamini and Hochberg,, 1995) the Abalone results - only significant at the 10% level with standard testing - are not significant anymore. All other results remain unchanged [^2] Due to the small sample sizes, Benjamini-Hochberg is preferred to the conservative Bonferroni adjustments (Nakagawa,, 2004)..
Table 1: Comparison between Mambular and XGBoost. The left tables shows regression results with average MSE values over 5 folds. The right side shows (binary) classification results with average AUC values. Significantly better values at the 5% significance level are in green and marked bold. Significantly better values at the 10% significance level are underscored. Dataset details can be found in appendix A. $\uparrow$ depicts higher is better and vice-versa.
Mambular 0.018 0.452 0.167 0.628 0.035 0.132 0.027 0.927 0.928 0.856 0.795 0.917
XGB 0.019 0.506 0.171 0.528 0.036 0.119 0.024 0.928 0.929 0.845 0.774 0.922
$p$-value 0.0079 0.0870 0.4865 1.3e-07 0.6287 0.3991 0.1999 0.7883 0.7930 0.0192 0.0120 0.010

### v2

@@ Abstract
The analysis of tabular data has traditionally been dominated by gradient-boosted decision trees (GBDTs), known for their proficiency with mixed categorical and numerical features. However, recent deep learning innovations are challenging this dominance. This paper investigates the use of autoregressive state-space models for tabular data and compares their performance against established benchmark models. Additionally, we explore various adaptations of these models, including different pooling strategies, feature interaction mechanisms, and bi-directional processing techniques to understand their effectiveness for tabular data. Our findings indicate that interpreting features as a sequence and processing them and their interactions through structured state-space layers can lead to significant performance improvement. This research underscores the versatility of autoregressive models in tabular data analysis, positioning them as a promising alternative that could substantially enhance deep learning capabilities in this traditionally challenging area. The source code is available at https://github.com/basf/mamba-tabular.
@@ 1 Introduction
Gradient-boosted decision trees (GBDTs) have long been the dominant approach for analyzing tabular data, due to their ability to handle the typical mix of categorical and numerical features found in such datasets (Grinsztajn et al.,, 2022). In contrast, deep learning models have historically faced challenges with tabular data, often struggling to outperform GBDTs. The complexity and diversity of tabular data, including issues like missing values, varied feature types, and the need for extensive preprocessing, have made it difficult for deep learning to match the performance of GBDTs (Borisov et al.,, 2022). However, recent advancements in deep learning are gradually challenging this paradigm by introducing innovative architectures that leverage advanced mechanisms to capture complex feature dependencies, promising significant improvements (Popov et al.,, 2019; Hollmann et al.,, 2022; Gorishniy et al.,, 2021).
@@
One of the most effective advancements in tabular deep learning is the application of attention mechanisms in models like TabTransformer (Huang et al.,, 2020), FT-Transformer (Gorishniy et al.,, 2021) and many more (Wang and Sun,, 2022; Thielmann et al., 2024b, ; Arik and Pfister,, 2021). These models leverage the attention mechanism to capture dependencies between features, offering a significant improvement over traditional approaches. FT-Transformers, in particular, have demonstrated robust performance across various tabular datasets, often surpassing the accuracy of GBDTs (McElfresh et al.,, 2024). Additionally, more traditional models like Multi-Layer Perceptrons (MLPs) and ResNets have demonstrated improvements when well-designed and when the data undergoes thorough preprocessing (Gorishniy et al.,, 2021, 2022). These models have benefited especially from innovations in advanced preprocessing methods that make them more competitive.
@@
More recently, the Mamba architecture (Gu and Dao,, 2023) has shown promising results in textual problems. Tasks previously dominated by Transformer architectures, such as DNA modeling and language modeling, have seen improvements with the application of Mamba models (Gu and Dao,, 2023; Schiff et al.,, 2024; Zhao et al.,, 2024). Several adaptations have demonstrated its versatility, such as Vision Mamba for image classification (Xu et al.,, 2024), video analysis (Yang et al.,, 2024; Yue and Li,, 2024) and point cloud analysis (Zhang et al.,, 2024; Liu et al.,, 2024). Furthermore, the architecture has been adapted for time series problems, with notable successes reported by Patro and Agneeswaran, (2024), Wang et al., (2024) and Ahamed and Cheng, 2024b . Mamba has also been integrated into graph learning (Behrouz and Hashemi,, 2024) and imitation learning (Correia and Alexandre,, 2024). Further advancements have improved the language model, for example, by incorporating attention (Lieber et al.,, 2024), Mixture of Experts (Pióro et al.,, 2024) or bi-directional sequence processing (Liang et al.,, 2024).
@@
These advancements underscore Mamba’s broad applicability, making it a powerful and flexible architecture for diverse tasks and data types. Similarly to the transformer architecture, the question arises whether the Mamba architecture can also be leveraged for tabular problems, and this study is focused on addressing this question.
@@
    We introduce a tabular adaptation of Mamba, subsequently called Mambular, demonstrating the potential of sequential models in addressing tabular problems.
@@
    We conduct extensive benchmarking of Mambular against several competitive neural and tree-based methods, illustrating that a standard Mambular model performs on par with or better than tree-based models across a wide range of datasets.
@@
    We examine the impact of bi-directional processing and feature interaction layers on Mambular’s performance, and compare several pooling methods.
@@
    Finally, we carry out an in-depth analysis of the models autoregressive structure, investigating the implications of feature orderings in a sequential tabular model.
[…]
@@ 2 Methodology
For a tabular problem, let $\mathcal{D}=\{(\bm{x}^{(i)},y^{(i)})\}_{i=1}^{n}$ be the training dataset of size $n$ and let $y$ denote the target variable that can be arbitrarily distributed. Each input $\bm{x}=(x_{1},x_{2},\dots,x_{J})$ contains $J$ features (variables). Categorical and numerical features are distinguished such that $\bm{x}\equiv(\bm{x}_{cat},\bm{x}_{num})$, with the complete feature vector denoted as $\bm{x}$. Further, let $x_{j(cat)}^{(i)}$ denote the $j$-th categorical feature of the $i$-th observation, and hence $x_{j(num)}^{(i)}$ denote the $j$-th numerical feature of the $i$-th observation.
Following standard tabular transformer architectures, the categorical features are first encoded and embedded. In contrast to classical language models, each categorical feature has its own, distinct vocabulary to avoid problems with binary or integer encoded variables. Including $<$unk$>$ tokens additionally allows to easily deal with unknown or missing categorical values during training or inference.
Numerical features are mapped to the embedding space via a simple linear layer. However, since a single linear layer does not add information beyond a linear transformation, Periodic Linear Encodings, as introduced by Gorishniy et al., (2022) are used for all numerical features. Thus, each numerical feature is encoded before being passed through the linear layer for rescaling. Simple decision trees are used for detecting the bin boundaries, $b_{t}$, and depending on the task, either classification or regression is employed for the target-dependent encoding function $h_{j}(\bm{x}_{j(num)},y)$. Let $b_{t}$ denote the decision boundaries from the decision trees. The encoding function is given in Eq. 1.
@@
Subsequently, the embeddings are passed jointly through a stack of Mamba layers. These include one-dimensional convolutional layers to account for invariance of feature ordering in the pseudo-sequence as well as a state-space (SSM) model (Gu et al.,, 2021; Hamilton,, 1994) . The feature matrix before being passed through the SSM model has a shape of (Batch Size) $\times$ J $\times$ (Embedding Dimension), later referenced as $N\times J\times d$. Importantly, the sequence length in a tabular context refers to the number of variables, and hence the second dimension, $J$, corresponds to the number of features rather than to the length of, e.g., a document.
The convolution operation along the sequence length $J$ and with Kernel $K$ is expressed as:
$\mathbf{Z}_{\text{conv}}^{(n,d)}(j)$ $=\sum_{m=0}^{K-1}\mathbf{Z}^{(n,d)}[j+m]\cdot\mathbf{k}^{(d)}(m),$
$\forall n$$\in\{1,\dots,N\},$
$\forall d$$\in\{1,\dots,d\},$
$\forall j$$\in\{1,\dots,J-K+1\}.$
where $\mathbf{Z}_{\text{conv}}^{(n,d)}(j)$ is the $j$-th element of the convolved sequence for batch $n$ and feature channel $d$. $\mathbf{Z}^{(n,d)}[j+m]$ is the $[j+m]$-th element of the input sequence $\mathbf{Z}$ for batch $n$ and feature channel $d$, and $K$ describes the kernel size. Summing over the elements of the kernel, indexed by $m$, accounts for the variable position in the pseudo-sequence. Thus, setting the kernel size equivalent to the number of variables would make the sequence invariant positional permutations. The resulting output tensor retains the same shape as the input, since padding is set to the kernel size -1.
After the convolution, given the matrices:
$\mathbf{A}\in\mathbb{R}^{1\times 1\times d\times\delta},\quad\mathbf{B}\in\mathbb{R}^{N\times J\times 1\times\delta},\quad\Delta\in\mathbb{R}^{N\times J\times d\times 1}$
$\bar{\mathbf{z}}\in\mathbb{R}^{N\times J\times d\times 1},$
@@
$\mathbf{h}_{j}=\exp\left(\Delta\odot_{3}\mathbf{A}\right)_{:,j,:,:}\odot_{1,2,3}\mathbf{h}_{j-1}+$ (2)
$\left(\left(\Delta\odot_{1,2}\mathbf{B}\right)\odot_{1,2,3}\bar{\mathbf{z}}\right)_{:,j,:,:}.$
@@
In contrast to FT-Transformer (Gorishniy et al.,, 2021) and TabTransformer (Huang et al.,, 2020) Mambular truly iterates through all variables as if they are a sequence; hence, feature interactions are detected sequentially. The effect of feature position in a sequence, and the impact of the convolution kernel size is analyzed with respect to performance in section 4. Furthermore, it should be noted that in contrast to TabPFN (Hollmann et al.,, 2022), Mambular does not transpose dimensions and iterates over observations. Hence, training on large datasets is possible and it can scale well to any training data size, just as Mamba (Gu and Dao,, 2023) does.
@@
Before passing $\tilde{\mathbf{x}}_{\text{final}}$ to the final task-specific model head, pooling is applied along the feature axis. Various strategies are analyzed in Section 4, with average pooling used as the default.
@@
The model is trained end-to-end by minimizing the task-specific loss, e.g., mean squared error for regression or categorical cross entropy for classification tasks. An overview of a forward pass in the model is given in Figure 2.
MambAttention
To accurately compare the introduced architecture with Transformer-based architectures, we also evaluate it against a model architecture, subsequently referred to as MambAttention, which is a simple combination of the proposed model and the FT-Transformer. This architecture loosely follows the modeling approach from Lieber et al., (2024). The input sequence is embedded as described previously, and the matrices $\mathbf{A},\mathbf{B},\Delta,\bar{\mathbf{z}}$ are constructed identically. However, after the first Mamba layer, a multi-head attention layer follows, similar to Gorishniy et al., (2021). The Mamba output representation, $\tilde{\mathbf{x}}_{\text{final}}\in\mathbb{R}^{N\times J\times d}$, is passed to the attention blocks in the same way as any embedded sequence representation. Following Huang et al., (2020), no positional embeddings are used, and, in contrast to Gorishniy et al., (2021), no [cls] token is introduced to maintain consistent dimensions across all model blocks. Subsequently, Mamba and attention blocks are used in an alternating fashion, with both the first and last block always being a Mamba block. The final model representation is obtained similarly to the Mambular model, by pooling over the output representation sequence.
@@ 3 Experiments
Figure 4: Critical difference diagram for all models on the benchmark datasets reported in Table 5. The average ranks across tasks are shown in brackets next to each model. Horizontal lines indicate groups of models with no statistically significant differences in performance. Mambular achieves the second-best average rank, with no significant differences at the 5% level to the best performing model, CatBoost. Notably, the top four models do not exhibit statistically significant differences from one another. The critical differences are computed using the Conover-Friedman test (Pereira et al.,, 2015), as both average ranks and performance metrics across all datasets are available.
Figure 5: Critical difference diagram for best performing models on additional results reported in Table 6. The average ranks across tasks are shown in brackets next to each model. Horizontal lines indicate groups of models with no statistically significant differences in performance. Mambular and CatBoost achieve the best average (identical) rank and significantly outperform the other models. The critical differences are computed using the Conover-Friedman test (Pereira et al.,, 2015), as both average ranks and performance metrics across all datasets are available.
The presented autoregressive architectures are evaluated against a range of top-performing models (McElfresh et al.,, 2024) across multiple datasets (Supplementary Table 7). These models include FT-Transformer (Gorishniy et al.,, 2021), TabTransformer (Huang et al.,, 2020), XGBoost (Grinsztajn et al.,, 2022; McElfresh et al.,, 2024), LightGBM (Ke et al.,, 2017), CatBoost (Prokhorenkova et al.,, 2018), NODE (Popov et al.,, 2019), a Random Forest, a baseline Multi-Layer Perceptron, and a ResNet. TabPFN (Hollmann et al.,, 2022) is excluded due to its unsuitability for larger datasets.
@@
A 5-fold cross-validation is conducted for all datasets, with average results and standard deviations reported. PLE encodings (Eq. 1) with a maximum number of bins equal to the model dimension are used for all neural models (128 for most models, including MLP and ResNet). All categorical features are integer-encoded. For regression tasks, targets are normalized. Mean Squared Error (MSE) and Area Under the Curve (AUC) metrics are reported for regression and classification tasks respectively. TabTransformer, FT-Transformer, Mambular and MambAttention all employ identical architectures for embeddings and task-specific heads, which includes a single output layer without activation function or dropout. The [cls] token embedding is utilized for final prediction in the FT-Transformer as it has been shown to enhance performance (Thielmann et al., 2024b, ; Gorishniy et al.,, 2021).
@@
All neural models share several parameters: a starting learning rate of 1e-04, weight decay of 1e-06, an early stopping patience of 15 epochs with respect to the validation loss, a maximum of 200 epochs for training, and learning rate decay with a factor of 0.1 with a patience of 10 epochs with respect to the validation loss. A universal batch size of 128 is used, and the best model with respect to the validation loss is returned for testing. TabTransformer, FT-Transformer, Mambular and MambAttention use the same embedding functions. For the benchmarks, a basic Mambular architecture is employed, using average pooling, no feature interaction layer, and no bi-directional processing. The columns/sequence are always sorted with numerical features first, followed by categorical features. Within these two groups, the features are sorted as they were originally provided in the dataset from the UCI Machine Learning Repository. A small kernel size of 4 in the convolutional layer is used based on the default Mamba architecture. The impact of variable positioning (with respect to the kernel size) on sequential processing is analyzed in section 4. Details on the used datasets and preprocessing can be found in Appendix B. Details on the model architectures and hyperparameters can be found in Appendix F.